# On the Global Optimality of Linear Policies for Sinkhorn Distributionally Robust Linear Quadratic Control

### v1

@@ IV-C Construction of an upper bound for $p^{\star}$
To derive an upper bound for $p^{\star}$, we restrict our attention to linear policies in $\mathcal{U}_{\eta}$ and appropriately enlarge the ambiguity set $\mathcal{S}$. This construction relies on the following result, which provides a lower bound to the Sinkhorn discrepancy between two distribution when one of them is Gaussian.
@@
Let $\mathbb{P}\sim\mathcal{N}(0,\Sigma_{1})$, $\Sigma_{1}\in\mathbb{S}_{+}^{d}$, and $\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d})$ be a distribution with covariance $\Sigma_{2}\in\mathbb{S}_{+}^{d}$. Then, it holds that $W_{\epsilon}(\mathbb{P},\mathbb{Q})\geq G_{\epsilon}(\Sigma_{1},\Sigma_{2})$.
  •   Proof:
@@
    and, using the relationships in [23, Proposition 2.1], that
@@
For any regularization parameter $\epsilon\geq 0$ and radius $\rho\geq 0$, it holds that the Sinkhorn ambiguity set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho\}$ is always contained in the entropy-regularized Gelbrich set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):\mathbb{E}_{\mathbb{Q}}[zz^{\top}]=\Sigma_{2}\in\mathbb{S}^{d}_{+},\ G_{\epsilon}(\Sigma_{1},\Sigma_{2})\leq\rho\}$.
  •   Proof:
    By Proposition 3, we have that $W_{\epsilon}(\mathbb{P},\mathbb{Q})\geq G_{\epsilon}(\Sigma_{1},\Sigma_{2})$. Hence, if $W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho$, then $G_{\epsilon}(\Sigma_{1},\Sigma_{2})\leq\rho$, which completes the proof. ∎
By restricting to liner policies, we finally construct an upper bound for (6) as
@@
As we enlarged nature’s subproblem feasible set and at the same time shrank the possible control laws that the designer can select, we have that $\overline{p}^{\star}\geq p^{\star}$. In the next proposition we rewrite (13) as a finite-dimensional optimization problem.
@@
  •   Proof:
    The proof follows the same steps as the proof of [14, Proposition 3.4]. In the same way as in Proposition 2, we first substitute $\bf{x=Hu+Gw}$ and $\bf{u=U(Dw+v)+q}$ in the objective of (13). Then, for any $\mathbb{P}\in\mathcal{G}$, we rewrite the previous expectation as a trace in terms of the covariance matrices $\mathbf{W}=\mathbb{E}_{\mathbb{P}}[\bf ww^{\top}]$ and $\mathbf{V}=\mathbb{E}_{\mathbb{P}}[\bf vv^{\top}]$, and conclude by replacing $\mathcal{G}$ with (10) and (11). ∎
@@ IV-D Existence of a Nash equilibrium
Given $\hat{\Sigma}\in\mathbb{S}^{d}_{+}$, $\rho\geq 0$ and $\epsilon\geq 0$ and finite, the set $\mathcal{D}=\{M\in\mathbb{S}^{d}_{+}:G_{\epsilon}(\hat{\Sigma},M)\leq\rho\}$ is convex and compact.
  •   Proof:
@@
    and $(\heartsuit)$ is convex because of [24, Proposition 6], while the trace is linear (hence convex) and the negative log-determinant is convex. This implies that $\mathcal{D}$ is convex because it is the level set of a convex function [29, Proposition 2.7].
    Compactness: we want to show that $\mathcal{D}$ is closed and bounded. To this end, let the function $f:\mathbb{S}^{d}_{+}\rightarrow\mathbb{R}$ be defined as
@@
    This function involves affine transformations of $M$ along with continuous transformations on $\mathbb{S}^{d}_{+}$ such as matrix square-root, trace and log-determinant. Hence, $f(\cdot)$ is continuous and the set $\mathcal{D}=\{M\in\mathbb{S}^{d}_{+}:f(M)\leq\tilde{\rho}\}$ with $\tilde{\rho}=\rho-\operatorname{Tr}(\hat{\Sigma})+\frac{\epsilon}{2}(\log(\frac{\epsilon}{2})^{d}-\log|\Sigma|)$ is closed because it is the lower level set of a continuous function [29, Theorem 1.6].
@@
    As $\lambda_{\max}(M)\rightarrow+\infty$, the linear terms in $\lambda_{\max}(M)$ dominate the other ones. Consequently, the RHS is unbounded when choosing $M$ such that $\lambda_{\max}(M)=+\infty$. Therefore, since $\tilde{\rho}$ is finite when so is $\epsilon$, we contradict the fact that any $M\in\mathcal{D}$ satisfies $f(M)\leq\tilde{\rho}$. Hence, $\mathcal{D}$ is bounded. ∎
Lemma 1 is key, as it enables the use of Sion’s minimax theorem [30, Theorem 3.4] to prove the existence of a Nash equilibrium for (4). This is formally stated in the next theorem.
@@
  •   Proof:
    By Lemma 1 the sets $\mathcal{G}_{W}$ and $\mathcal{G}_{V}$ are compact and convex. The trace is linear, hence the objective function is concave in $\mathbf{W}$ and $\mathbf{V}$. Moreover, since $\mathbf{Q},\mathbf{R},\mathbf{M}\succeq 0$, it is convex in $\mathbf{U}$ and $\mathbf{q}$. Therefore, we can apply Sion’s minimax theorem [30, Theorem 3.4] to show that strong duality holds, proving the first point of the theorem. By strong duality, the chain of inequalities $\underline{d}^{\star}\leq d^{\star}\leq p^{\star}\leq\overline{p}^{\star}$ collapses to equalities proving the second point. The equality $p^{\star}=\overline{p}^{\star}$ implies that (6) is solved by a linear causal policy of the purified observations. However, as pointed out in Section IV-A, any causal controller that is linear in the purified outputs $\bm{\eta}$ can be also expressed as a causal linear feedback in the measurements $\bf y$. This proves the third point. We finish the proof noticing that the last point of the theorem follows from the identity $\underline{d}^{\star}=d^{\star}$. ∎
We conclude this section by observing that, when $\epsilon\rightarrow\infty$, the set $\mathcal{S}$ either becomes empty or reduces to the singleton $\nu$ depending on whether $\operatorname{Tr}(\hat{\Sigma})+\operatorname{Tr}(\Sigma)$ exceeds $\rho$ or not, see [21, Proposition 1]. In particular, when $\operatorname{Tr}(\hat{\Sigma})+\operatorname{Tr}(\Sigma)\leq\rho$, we still retain global optimality of linear policies as $\mathcal{S}=\{\nu\}$, and $\nu$ is Gaussian. Instead, when $\operatorname{Tr}(\hat{\Sigma})+\operatorname{Tr}(\Sigma)>\rho$, the Sinkhorn DR LQG problem (4) becomes unfeasible.
@@ V Numerical Experiments
In this section, we present numerical simulations to showcase the advantages of robustifying to distributional uncertainty. [^3] All our experiments were run on a M3 Pro CPU machine with 36GB RAM. All SDP problems were modeled in Matlab 2023a using Yalmip and solved with MOSEK. Our source code is publicly available at https://github.com/DecodEPFL/Optimality_Sinkhorn.git. For our experiments, we consider the open-loop unstable discrete-time linear dynamical system given by
$x_{t+1}=\begin{bmatrix}1.1&0.1\\ 0&1.1\end{bmatrix}x_{t}+u_{t}+w_{t},\quad y_{t}=x_{t}+v_{t}\,$,
with cost matrices $Q_{t}=I_{2}$ and $R_{t}=10^{-3}\cdot I_{2}$ at all times, control horizon $T=25$, and nominal covariance $\hat{\Sigma}=I$, that is, an identity matrix of appropriate dimensions.
In Fig. 1, we benchmark the classical LQG controller designed based on $\hat{\Sigma}$ against the Sinkhorn DR LQG policy obtained by solving (4) with radii $\rho_{x_{0}}=\rho_{w_{t}}=\rho_{v_{t}}=10^{3}$ and regularization parameter $\epsilon=1$. Specifically, we compare the performance of these controllers on $5000$ realizations of the exogenous disturbances drawn from the nominal Gaussian distribution (on the left) and from the respective nature’s optimal choice of distribution $\mathbb{P}$ within $\mathcal{S}$ (on the right). As expected, we observe that the Sinkhorn DR LQG policy incurs a slightly higher average cost when the true distribution corresponds to the nominal one. Conversely, when the noise distributions are selected adversarially within $\mathcal{S}$, we observe that the proposed Sinkhorn DR LQG policy achieves a lower average cost. These results validate our design and highlight fundamental tradeoff between nominal performance and distributional robustness.
Figure 1: Comparison between the control cost incurred by the nominal LQG controller and the proposed Sinkhorn DR LQG policy over 50005000 disturbance realizations drawn from the nominal distribution (on the left) and the respective worst-case distribution in the Sinkhorn ambiguity set (on the right). Dotted vertical lines represent theoretical mean values.
@@ VI Conclusion
In this work, we studied a DR generalization of classical LQG control where the noise distributions are unknown and belong to entropy-regularized Wasserstein or Sinkhorn ambiguity sets. We proved that, despite robustifying the objective in the probability space, nature’s distributions retain a Gaussian form and hence linear policies remain globally optimal. We validated the effectiveness of our Sinkhorn DR LQG policy through numerical simulations showing its improved robustness compared to classical LQG design.
@@ Appendix A Appendix
We report below expressions for the stacked cost matrices $\mathbf{Q}\in\mathbb{S}^{d(T+1)}$ and $\mathbf{R}\in\mathbb{S}_{+}^{mT}$, and for the matrices $\mathbf{C}\in\mathbb{R}^{pT\times d(T+1)}$, $\mathbf{G}\in\mathbb{R}^{d(T+1)\times d(T+1)}$, and $\mathbf{H}\in\mathbb{R}^{d(T+1)\times mT}$ encoding the system linear dynamics and measurement equations. Specifically, we have $\mathbf{Q}=\operatorname{blkdiag}(Q_{0},\dots,Q_{T})$, $\mathbf{R}=\operatorname{blkdiag}(R_{0},\dots,R_{T-1})$ and

### v2

@@ IV-C Construction of an upper bound for $p^{\star}$
To derive an upper bound for $p^{\star}$, we restrict our attention to linear policies in $\mathcal{U}_{\eta}$ and appropriately enlarge the ambiguity set $\mathcal{S}$. This construction relies on the following result, which provides a lower bound to the Sinkhorn discrepancy between two distributions when one of them is Gaussian.
@@
Let $\mathbb{P}\sim\mathcal{N}(0,\Sigma_{1})$, $\Sigma_{1}\in\mathbb{S}_{++}^{d}$, and $\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d})$ be a distribution with covariance $\Sigma_{2}\in\mathbb{S}_{++}^{d}$. Then, it holds that $W_{\epsilon}(\mathbb{P},\mathbb{Q})\geq G_{\epsilon}(\Sigma_{1},\Sigma_{2})$.
@@
and, using the relationships in [7, Proposition 2.1], that
@@
For any regularization parameter $\epsilon\geq 0$ and radius $\rho\geq 0$, it holds that the Sinkhorn ambiguity set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho\}$ is always contained in the entropy-regularized Gelbrich set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):\mathbb{E}_{\mathbb{Q}}[zz^{\top}]=\Sigma_{2}\in\mathbb{S}^{d}_{++},\ G_{\epsilon}(\Sigma_{1},\Sigma_{2})\leq\rho\}$.
By Proposition 3, we have that $W_{\epsilon}(\mathbb{P},\mathbb{Q})\geq G_{\epsilon}(\Sigma_{1},\Sigma_{2})$. Hence, if $W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho$, then $G_{\epsilon}(\Sigma_{1},\Sigma_{2})\leq\rho$, which completes the proof. ∎ By restricting to linear policies, we finally construct an upper bound for (7) as
@@
As we enlarged nature’s subproblem feasible set, and at the same time shrank the possible control laws that the designer can select, we have that $\overline{p}^{\star}\geq p^{\star}$. In the next proposition, we rewrite (14) as a finite-dimensional optimization problem.
@@
The proof follows the same steps as the proof of [28, Proposition 3.4]. In the same way as in Proposition 2, we first substitute $\bf{x=Hu+Gw}$ and $\bf{u=U(Dw+v)+q}$ in the objective of (14). Then, for any $\mathbb{P}\in\mathcal{G}$, we rewrite the previous expectation as a trace in terms of the covariance matrices $\mathbf{W}=\mathbb{E}_{\mathbb{P}}[\bf ww^{\top}]$ and $\mathbf{V}=\mathbb{E}_{\mathbb{P}}[\bf vv^{\top}]$, and conclude by replacing $\mathcal{G}$ with (11) and (12). ∎
@@ IV-D Existence of a Nash equilibrium
Given $\hat{\Sigma}\in\mathbb{S}^{d}_{++}$, $\rho\geq 0$ and $\epsilon\geq 0$ and finite, the set $\mathcal{D}=\{M\in\mathbb{S}^{d}_{++}:G_{\epsilon}(\hat{\Sigma},M)\leq\rho\}$ is convex and compact.
@@
and $(\heartsuit)$ is convex because of [18, Proposition 6], while the trace is linear (hence convex) and the negative log-determinant is convex. This implies that $\mathcal{D}$ is convex because it is the level set of a convex function [26, Proposition 2.7].
Compactness: we want to show that $\mathcal{D}$ is closed and bounded. To this end, let the function $f:\mathbb{S}^{d}_{++}\rightarrow\mathbb{R}$ be defined as
@@
This function involves affine transformations of $M$ along with continuous transformations on $\mathbb{S}^{d}_{++}$ such as matrix square-root, trace and log-determinant. Hence, $f(\cdot)$ is continuous and the set $\mathcal{D}=\{M\in\mathbb{S}^{d}_{++}:f(M)\leq\tilde{\rho}\}$ with $\tilde{\rho}=\rho-\operatorname{Tr}(\hat{\Sigma})+\frac{\epsilon}{2}(\log(\frac{\epsilon}{2})^{d}-\log|\Sigma|)$ is closed because it is the lower level set of a continuous function [26, Theorem 1.6].
@@
As $\lambda_{\max}(M)\rightarrow+\infty$, the linear terms in $\lambda_{\max}(M)$ dominate the other ones. Consequently, the RHS is unbounded when choosing $M$ such that $\lambda_{\max}(M)=+\infty$. Therefore, since $\tilde{\rho}$ is finite when so is $\epsilon$, we contradict the fact that any $M\in\mathcal{D}$ satisfies $f(M)\leq\tilde{\rho}$. Hence, $\mathcal{D}$ is bounded. ∎ Lemma 1 is key, as it enables the use of Sion’s minimax theorem [27, Theorem 3.4] to prove the existence of a Nash equilibrium for (5). This is formally stated in the next theorem.
@@
By Lemma 1 the sets $\mathcal{G}_{W}$ and $\mathcal{G}_{V}$ are compact and convex. The trace is linear, hence the objective function is concave in $\mathbf{W}$ and $\mathbf{V}$. Moreover, since $\mathbf{Q},\mathbf{R},\mathbf{M}\succeq 0$, it is convex in $\mathbf{U}$ and $\mathbf{q}$. Therefore, we can apply Sion’s minimax theorem [27, Theorem 3.4] to show that strong duality holds, proving the first point of the theorem. By strong duality, the chain of inequalities $\underline{d}^{\star}\leq d^{\star}\leq p^{\star}\leq\overline{p}^{\star}$ collapses to equalities proving the second point. The equality $p^{\star}=\overline{p}^{\star}$ implies that (7) is solved by a linear causal policy of the purified observations. However, as pointed out in Section IV-A, any causal controller that is linear in the purified outputs $\bm{\eta}$ can be also expressed as a causal linear feedback in the measurements $\bf y$. This proves the third point. We finish the proof noticing that the last point of the theorem follows from the identity $\underline{d}^{\star}=d^{\star}$. ∎ Theorem 1 implies that (5) is solved by an LQG with worst-case covariance matrices given by (10). We report a reformulation of (10) as finite-dimensional conic program in Proposition 5 in the Appendix. We conclude this section by observing that, when $\epsilon\rightarrow\infty$, the set $\mathcal{S}$ either becomes empty or reduces to the singleton $\nu$ depending on whether $\operatorname{Tr}(\hat{\Sigma})+\operatorname{Tr}(\Sigma)$ exceeds $\rho$ or not, see [6, Proposition 1]. In particular, when $\operatorname{Tr}(\hat{\Sigma})+\operatorname{Tr}(\Sigma)\leq\rho$, we still retain global optimality of linear policies as $\mathcal{S}=\{\nu\}$, and $\nu$ is Gaussian. Instead, when $\operatorname{Tr}(\hat{\Sigma})+\operatorname{Tr}(\Sigma)>\rho$, the Sinkhorn DR LQG problem (5) becomes infeasible.
@@ V Numerical Experiments
In this section, we present numerical simulations to showcase the advantages of robustifying against distributional uncertainty. [^4] All our experiments were run on an M3 Pro CPU machine with 36GB RAM. All SDP problems were modeled in Matlab 2023a using Yalmip and solved with MOSEK. Our source code is publicly available at https://github.com/DecodEPFL/Optimality_Sinkhorn.git. For our experiments, we consider the open-loop unstable discrete-time linear dynamical system given by
$x_{t+1}=\begin{bmatrix}1.1&0.1\\ 0&1.1\end{bmatrix}x_{t}+\begin{bmatrix}1\\ 1\end{bmatrix}u_{t}+w_{t},\quad y_{t}=x_{t}+v_{t}\,$,
with cost matrices $Q_{t}=I_{2}$ and $R_{t}=1$ at all times, control horizon $T=25$, and nominal covariances $\hat{X}_{0}=I_{2},\hat{W}_{t}=I_{2},\hat{V}_{t}=0.01I_{2},\ \forall t\in[T]$. The reference covariance was picked $\Sigma=I_{2}$.
In Fig. 1, we benchmark the classical LQG controller designed based on the nominal covariances against the Sinkhorn DR LQG policy obtained by solving (5) using the conic program in Proposition 5 with radii $\rho_{x_{0}}=20,\rho_{w_{t}}=0.2,\rho_{v_{t}}=20$ and regularization parameter $\epsilon=0.01$. Specifically, we compare the performance of these controllers on $5000$ realizations of the exogenous disturbances drawn from the nominal Gaussian distribution (on the left) and from the respective nature’s adversarial choice of distribution $\mathbb{P}$ within $\mathcal{S}$ (on the right). As expected, we observe that the Sinkhorn DR LQG policy incurs a slightly higher average cost when the true distribution corresponds to the nominal one. Conversely, when the noise distributions are selected adversarially within $\mathcal{S}$, we observe that the proposed Sinkhorn DR LQG policy achieves a lower average cost. These results validate our design and highlight fundamental tradeoff between nominal performance and distributional robustness.
Figure 1: Comparison between the control cost incurred by the nominal LQG controller (red histograms, in the background) and the proposed Sinkhorn DR LQG policy (green histograms, in the foreground) over 50005000 disturbance realizations drawn from the nominal distribution (on the left) and the respective worst-case distribution in the Sinkhorn ambiguity set (on the right). Dotted vertical lines represent theoretical mean values.
@@ VI Conclusion
In this work, we studied a DR generalization of classical LQG control where the noise distributions are unknown and belong to entropy-regularized Wasserstein or Sinkhorn ambiguity sets. We proved that, despite robustifying the objective in the probability space, nature’s distributions retain a Gaussian form and hence linear policies remain globally optimal. We validated the effectiveness of our Sinkhorn DR LQG policy through numerical simulations showing improved robustness compared to classical LQG design.
@@ Appendix A Appendix
We report below expressions for the stacked cost matrices $\mathbf{Q}\in\mathbb{S}_{+}^{d(T+1)}$ and $\mathbf{R}\in\mathbb{S}_{++}^{mT}$, and for the matrices $\mathbf{C}\in\mathbb{R}^{pT\times d(T+1)}$, $\mathbf{G}\in\mathbb{R}^{d(T+1)\times d(T+1)}$, and $\mathbf{H}\in\mathbb{R}^{d(T+1)\times mT}$ encoding the system linear dynamics and measurement equations. Specifically, we have $\mathbf{Q}=\operatorname{blkdiag}(Q_{0},\dots,Q_{T})$, $\mathbf{R}=\operatorname{blkdiag}(R_{0},\dots,R_{T-1})$ and
@@
Proposition 5:
The optimization problem (10) is equivalent to the conic program
$\max\$$\operatorname{Tr}(\mathbf{G^{\top}QGW})-\operatorname{Tr}\big(\mathbf{M}^{-1}F\big)$
s.t. $\Gamma\in\mathcal{M},\ F\in\mathbb{S}_{+}^{mT},\mathbf{W}\in\mathbb{S}_{++}^{d(T+1)},\ \mathbf{V}\in\mathbb{S}_{++}^{pT},$
$E_{x_{0}}\in\mathbb{S}_{+}^{d},\ E_{w_{t}}\in\mathbb{S}_{+}^{d},\ E_{v_{t}}\in\mathbb{S}_{+}^{p},\ \forall t\in[T],$
$\begin{aligned} &\operatorname{Tr}\!\left(\!W_{0}\!-\!2E_{x_{0}}\!+\!\frac{\epsilon}{2}\Sigma^{-1}W_{0}\!\right)\!-\!\frac{\epsilon}{2}\log\left|E_{x_{0}}\!-\frac{\epsilon}{4}I\right|\leq\rho_{x_{0}}\\ &-\operatorname{Tr}(\hat{X}_{0})+\frac{\epsilon}{2}\left(d\log\left(\frac{\epsilon}{2}\right)-\log|\hat{X}_{0}\Sigma|\right),\end{aligned}$
$\begin{aligned} &\operatorname{Tr}\left(\!W_{t+1}\!\!-\!2E_{w_{t}}\!\!+\!\frac{\epsilon}{2}\Sigma^{-1}W_{t+1}\!\right)\!\!-\!\frac{\epsilon}{2}\!\log\!\left|E_{w_{t}}\!\!-\!\frac{\epsilon}{4}I\right|\!\!\leq\!\rho_{w_{t}}\\ &-\operatorname{Tr}(\hat{W}_{t})+\frac{\epsilon}{2}\left(d\log\left(\frac{\epsilon}{2}\right)-\log|\hat{W}_{t}\Sigma|\right),\ \forall t\in[T]\end{aligned}$
$\begin{aligned} &\operatorname{Tr}\left(V_{t}\!-\!2E_{v_{t}}\!+\!\frac{\epsilon}{2}\Sigma^{-1}V_{t}\right)\!-\!\frac{\epsilon}{2}\log\left|E_{v_{t}}\!-\!\frac{\epsilon}{4}I\right|\leq\rho_{v_{t}}\\ &-\operatorname{Tr}(\hat{V}_{t})+\frac{\epsilon}{2}\left(d\log\left(\frac{\epsilon}{2}\right)-\log|\hat{V}_{t}\Sigma|\right),\ \forall t\in[T]\end{aligned}$
$\begin{bmatrix}\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}&E_{x_{0}}\\ E_{x_{0}}&I_{n}\end{bmatrix}\succeq 0,$
$\begin{bmatrix}\hat{W}_{t}^{1/2}W_{t+1}\hat{W}_{t}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}&E_{w_{t}}\\ E_{w_{t}}&I_{n}\end{bmatrix}\succeq 0,\ \forall t\in[T],$
$\begin{bmatrix}\hat{V}_{t}^{1/2}V_{t}\hat{V}_{t}^{1/2}+\frac{\epsilon^{2}}{16}I_{p}&E_{v_{t}}\\ E_{v_{t}}&I_{p}\end{bmatrix}\succeq 0,\ \forall t\in[T],$
$\begin{bmatrix}F&\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2\\[1.0pt] (\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2)^{\top}&\mathbf{DWD^{\top}+V}\end{bmatrix}\!\!\succeq\!0\,.$
Here, $\mathcal{M}$ denotes the set of all strictly upper block triangular matrices of the form
$\Gamma=\begin{bmatrix}0&\Gamma_{1,2}&\Gamma_{1,3}&\cdots&\Gamma_{1,T}\\ &0&\Gamma_{2,3}&\cdots&\Gamma_{2,T}\\ &&\ddots&\ddots&\vdots\\ &&&0&\Gamma_{T-1,T}\\ &&&&0\end{bmatrix}\in\mathbb{R}^{Tm\times Tp}\,$,
where $\Gamma_{t,s}\in\mathbb{R}^{m\times p}$ for every $t,s\in\mathbb{Z}$ with $1\leq t<s\leq T$.
The proof relies on dualizing the inner minimization problem in (10). Strong duality holds because the primal problem is feasible and involves only equality constraints, hence any feasible point is in fact a Slater point.
In the following we use $\Gamma\in\mathcal{M}$ to denote the Lagrange multiplier of the constraint $\bf U\in\mathcal{U}^{\text{lin}}$, which requires all blocks of $\bf U$ above the main diagonal to vanish. The Lagrangian function of the inner minimization problem in (10) is
$\mathcal{L}(\mathbf{q},\mathbf{U},\Gamma)=\operatorname{Tr}\left(\mathbf{F}_{1}\mathbf{W}+\mathbf{F}_{2}\mathbf{V}\right)+\mathbf{q}^{\top}\mathbf{M}\mathbf{q}+\operatorname{Tr}(\mathbf{U}\Gamma^{\top})$
Recall now that $\mathbf{R}\succ 0$ and $\mathbf{Q}\succeq 0$, and thus $\mathbf{M}\succ 0$. Consequently, $\mathcal{L}$ is minimized by $\mathbf{q}^{\star}=0$ for any fixed $\bf U$ and $\Gamma$. In addition, the partial gradient of $\mathcal{L}$ with respect to $\bf U$ is given by
$\frac{\partial\mathcal{L}}{\partial\bf U}=2\mathbf{MUDWD}^{\top}+2\mathbf{MUV}+2\mathbf{H^{\top}QGWD^{\top}}+\Gamma\,.$
Recall also that $\mathbf{V}\in\mathcal{G}_{V}$ is strictly positive, which implies that $\mathbf{DWD^{\top}}+\mathbf{V}\succ 0$ is invertible. As we already know that $\mathbf{M}\succ 0$ is invertible as well, $\mathcal{L}$ is minimized by
$\mathbf{U}^{\star}=-\mathbf{M}^{-1}(\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2)(\mathbf{DWD^{\top}+V})^{-1}$
for any fixed $\Gamma$.
Substituting both $\bf q^{\star}$ and $\bf U^{\star}$ into $\mathcal{L}$ yields the dual objective function
$g(\Gamma)=\mathcal{L}(\mathbf{q^{\star},U}^{\star},\Gamma)=\operatorname{Tr}(\mathbf{G^{\top}QGW})+$
[…]
The dual of the inner minimization problem in (10) is thus given by $\max_{\Gamma\in\mathcal{M}}g(\Gamma)$. To linearize the dual function $g(\Gamma)$, we introduce the auxiliary variable $F\in\mathbb{S}_{+}^{mT}$ subject to the matrix inequality
[…]
We can reformulate the previous inequality as an LMI using the Schur complement. Then, the dual problem rewrites as
$\max$ $\operatorname{Tr}(\mathbf{G^{\top}QGW})-\operatorname{Tr}\big(\mathbf{M}^{-1}F\big)$ (16)
s.t. $\Gamma\in\mathcal{M},\ F\in\mathbb{S}_{+}^{mT},$
$\!\!\begin{bmatrix}F&\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2\\[1.0pt] (\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2)^{\top}&\mathbf{DWD^{\top}+V}\end{bmatrix}\!\!\succeq\!\!0\,.$
Substituting the strong dual (16) in the inner minimization in (10), we obtain
$\max\ \$$\operatorname{Tr}(\mathbf{G^{\top}QGW})-\operatorname{Tr}\big(\mathbf{M}^{-1}F\big)$
s.t. $\Gamma\in\mathcal{M},\ F\in\mathbb{S}_{+}^{mT},\mathbf{W}\in\mathbb{S}_{++}^{d(T+1)},\ \mathbf{V}\in\mathbb{S}_{++}^{pT},$
$\!\!\begin{bmatrix}F&\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2\\[1.0pt] (\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2)^{\top}&\mathbf{DWD^{\top}+V}\end{bmatrix}\!\!\succeq\!0,$
$G_{\epsilon}(\hat{X}_{0},W_{0})\leq\rho_{x_{0}},$
$G_{\epsilon}(\hat{W}_{t+1},W_{t+1})\leq\rho_{w_{t}},\ G_{\epsilon}(\hat{V}_{t},V_{t})\leq\rho_{v_{t}}\ \forall t\in[T]\,.$
We conclude the proof by reformulating the entropy-regularized Gelbrich constraints. For simplicity, consider the constraint $G_{\epsilon}(\hat{X}_{0},W_{0})\leq\rho_{x_{0}}$; the reasoning applies mutatis mutandis to the others. Consider an auxiliary variable $E_{x_{0}}\in\mathbb{S}^{n}_{+}$ subject to the matrix inequality $E_{x_{0}}^{2}\preceq\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}$. This inequality can be recast as $E_{x_{0}}\preceq(\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n})^{\frac{1}{2}}$. With this auxiliary variable, we can equivalently express the nonlinear inequality constraint as
[…]
If the constraint is satisfied with $E_{x_{0}}$ it will be also with the original variable since the trace and the log-determinant are monotone operators. Finally, with a Schur complement argument the inequality $E_{x_{0}}^{2}\preceq\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}$ can be equivalently rewritten as
$\begin{bmatrix}\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}&E_{x_{0}}\\ E_{x_{0}}&I_{n}\end{bmatrix}\succeq 0\,.$